# Quantum metrology enhanced by effective time reversal

### v1

@@ Abstract
Quantum metrology involves the application of quantum resources to enhance measurements. Several communities have developed quantum-metrology strategies that leverage effective time reversals. These strategies, we posit, form four classes. First, echo metrology begins with a preparatory unitary and ends with that unitary’s time-reverse. The protocol amplifies the visibility of a small parameter to be sensed. Similarly, weak-value amplification enhances a weak coupling’s detectability. The technique exhibits counterintuitive properties captured by a retrocausal model. Using the third strategy, one simulates closed timelike curves, worldlines that loop back on themselves in time. The fourth strategy involves indefinite causal order, which characterises channels applied in a superposition of orderings. We review these four strategies, which we unify under the heading of time-reverse metrology. We also outline opportunities for this toolkit in quantum metrology; quantum information science; quantum foundations; atomic, molecular, and optical physics; and solid-state physics.
@@ I Introduction
    We can prepare a sensor with a unitary operation $\hat{V}$, couple the sensor to a target system, and then subject the sensor to $\hat{V}^{\dagger}$. The $\hat{V}$ and $\hat{V}^{\dagger}$ magnify the target’s impact on the sensor state. We call this strategy echo metrology.
@@
Echo Weak-value Time-loop ICO metrology
@@
Echo In $\leftarrow$, a unitary realises Some echo- In $\leftarrow$, we invert a
@@ II Technical introduction
Often, when performing parameter estimation, we infer the parameter $\alpha\in\mathbb{R}$ in a unitary $\hat{U}_{\alpha}\coloneqq e^{-i\alpha\hat{H}}$. The generator $\hat{H}$ is a Hermitian operator. Some target system produces $\hat{U}_{\alpha}\,$; example targets include an external field, as well as a physical system that couples to the probe. We aim to infer $\alpha$ with the least error possible, given our resources. [^2] In some time-loop metrology (Sec. V) and ICO metrology (Sec. VI), the inferred parameter characterises a nonunitary channel. We illustrate parameter estimation here with $\hat{U}_{\alpha}$ for simplicity and due to $U_{\alpha}$’s role in echo metrology (Sec. III), weak-value amplification (Sec. IV), some time-loop metrology, and some ICO metrology.
@@
III Echo metrology
Echo metrology involves a unitary $\hat{V}$ and the time-reverse $\hat{V}^{\dagger}$ thereof. Section III A motivates the protocol intuitively. In Sec. III B, we detail the protocol and prove that it achieves the information-bearing probe state’s QFI. Section III C illustrates the protocol with examples that involve squeezing. Section III D describes opportunities for future work.
Physicists have called many phenomena echoes [70]. During a Loschmidt echo, we evolve a system forward for a duration under a Hamiltonian $\hat{H}$, then backward for the same duration under a perturbed Hamiltonian $\hat{H}^{\prime}$. In nuclear-magnetic-resonance experiments, spin echos mitigate decoherence. This section’s echo protocol differs: during it, we prepare the probe unitarily, perturb the probe with the target, and reverse the unitary.
III A Intuition behind echo metrology
The following story helps motivate echo metrology, although without explaining its behaviour rigorously. Consider a classical chaotic system, such as a double pendulum. It undergoes the protocol sketched in Fig. 1: the system begins at some point in phase space, then evolves under its chaotic Hamiltonian. A small perturbation nudges the system’s phase-space point. Then, the system evolves backward in time under the Hamiltonian. The chaos magnifies the typical such perturbation, by definition.
Figure 1: Motivation for echo metrology: Sketch of a classical chaotic system’s trajectory through phase space. The system evolves forward under its Hamiltonian, undergoes a perturbation, and evolves backward under its Hamiltonian. The chaos magnifies the perturbation, rendering it easier to detect.
This protocol follows a pattern similar to that of echo metrology. The perturbation above plays the role of the $e^{-i\alpha\hat{H}}$ introduced in Sec. II. Forward and reverse evolutions enhance the signal’s detectability. References [102, 104, 1] suggested the analogy. However, echo metrology need not involve chaos; it relies on nontrivial, if not necessarily chaotic, evolution.
III B Introduction to, and analysis of, echo metrology
Below, we detail echo metrology. We then prove that it achieves the QFI of the information-bearing probe state. Finally, we review literature about echo metrology (that does not involve squeezing, covered in Sec. III C).
Echo metrology features the following setup and protocol. Consider a quantum probe whose Hilbert space contains a simple (unentangled, unsqueezed, etc.) state $\lvert 0\rangle$. A Hamiltonian $\hat{H}$ generates a unitary $e^{-i\alpha\hat{H}}$ parameterised by an $\alpha\in\mathbb{R}$ of magnitude $|\alpha|\ll 1$. The following protocol implements echo metrology:
@@ III B Introduction to, and analysis of, echo metrology
The final equality follows from step 1 of the protocol. This FI equals the QFI of $e^{-i\alpha\hat{H}}\,|\psi\rangle$ at $\mathcal{O}(\alpha)$, by Eq. (5). Thus, echo metrology achieves the QFI.
Previous works state or implicitly leverage this result. Reference [115] synopsises the argument above. In [33], a similar argument focusses on certain $\hat{V}$s and $\hat{H}$s. Other references address general echo metrology, arguing that it achieves the QFI of $e^{-i\alpha\hat{H}}\hat{V}\lvert 0\rangle$ [105, 176]. The review [3] illustrates echo metrology with examples. Researchers have also extended echo metrology to include ancilla systems [33].
Experiments have realised echo metrology. The protocol has been performed (in the absence of squeezing) with trapped ions [100, 101, 117], a Rydberg atom and an oscillator [133], and nitrogen-vacancy centres in diamond [147]. Many optical interferometers implement echo metrology, casting $V$ as a beamsplitter unitary [26, 145].
III C Echo metrology based on squeezing
In much echo metrology, the preparation involves squeezing. Squeezing is an operation that reduces one observable’s uncertainty at the expense of another observable’s [77]. We illustrate squeezing-based echo metrology with single-mode squeezing, spin squeezing, and two-mode squeezing.
First, we motivate single-mode echo metrology. Consider a bosonic mode – equivalently, a harmonic oscillator – associated to a position operator $\hat{Q}$ and a momentum operator $\hat{P}$. Suppose that a field imprints a parameter $\alpha$ on the oscillator’s state at a rate $g$, through the Hamiltonian $\hat{H}=g\alpha\hat{Q}$. This Hamiltonian kicks the oscillator’s momentum: $e^{i\hat{H}t}\,\hat{P}\,e^{-i\hat{H}t}=\hat{P}-g\alpha t\hat{\mathbbm{1}}$. The momentum’s displacement reveals $\alpha$, if we know $g$. Suppose we have prepared the oscillator in a state $\lvert\psi\rangle$. The momentum displacement’s expectation value forms an estimator:
@@ III C Echo metrology based on squeezing
Squeezing facilitates echo metrology for sensing $\alpha$. The probe begins in $\lvert 0\rangle$, undergoes squeezing by $e^{-ir\hat{H}_{\rm sq}(\varphi)}$, and evolves under $\hat{H}$. We then antisqueeze the sensor (evolve it under $e^{ir\hat{H}_{\rm sq}(\varphi)}$). Finally, we measure $\{\lvert 0\rangle\!\langle 0\rvert,\mathbbm{1}-\lvert 0\rangle\!\langle 0\rvert\}$. The unitaries condense into [^6] We apply the identity $e^{ir\hat{H}_{\rm sq}(\varphi{=}\pi/2)}\,\hat{Q}\,e^{-ir\hat{H}_{\rm sq}(\varphi{=}\pi/2)}=e^{r}\,\hat{Q}$.
@@
Not only a bosonic mode, but also a collection of spins, can undergo squeezing. [^7] The spins are typically qubits. However, we use the term spins for generality and for consistency with the literature. Various Hamiltonians enact spin squeezing, as discussed in App. A. Examples include the one-axis-twisting [94], the two-axis-twisting [94], the Lipkin-Meshkov-Glick [106], and general spin–spin interacting [19] Hamiltonians. Echo metrology offers two benefits when implemented with spin squeezing. First, an estimator’s variance can scale in Heisenberg fashion with the number of spins [40]. Second, the strategy offers robustness with respect to readout error [40, 124, 97]. Neutral-atom experiments have showcased these advantages [79, 37, 104].
An early squeezing-based echo-metrology proposal involves two-mode squeezing [179]. The protocol features two bosonic modes, associated to annihilation operators $\hat{a}$ and $\hat{b}$. Suppose that mode $a$ can undergo the phase shift $e^{i\alpha\hat{a}^{\dagger}\hat{a}}$. To sense $\alpha$, we prepare both modes in their vacuum states: $\lvert 0,0\rangle$. The unitary $\hat{S}(r)\coloneqq e^{r(\hat{a}^{\dagger}\hat{b}^{\dagger}-\hat{b}\hat{a})}$ introduces two-mode squeezing. Mode $a$ undergoes the phase shift. The unitary $\hat{S}^{-1}(r)$ unsqueezes the modes. Finally, we measure each mode’s photon number. This protocol achieves the QFI of $e^{i\alpha\hat{a}^{\dagger}\hat{a}}\,\hat{S}(r)\lvert 0,0\rangle$.
@@
III D Outlook for echo metrology
The echo-metrology template offers opportunities in and beyond the parameter estimation of Sec. II. Below, we present open questions about implementations. We then summarise possible applications to quantum information processing outside parameter estimation. Section V details how certain echo-metrology experiments simulate CTCs.
The echo-metrology protocol begins with a unitary $\hat{V}$ (step 1 in Sec. III B). What, generally, is the optimal form of $\hat{V}$? Consider any $\hat{V}$ that prepares an equal-weight superposition $(\lvert h_{\rm min}\rangle+e^{-i\phi}\lvert h_{\rm max}\rangle)/\sqrt{2}\,$, wherein $\phi\in[0,2\pi)$. Every such $\hat{V}$ can provide the QFI, maximised over $\hat{V}^{\prime}$, of $e^{-i\alpha\hat{H}}\hat{V}^{\prime}\lvert 0\rangle$ [Eq. (5)]. Yet implementing such $\hat{V}$s may pose challenges; for example, they may cost many computational gates. Furthermore, noise can affect the optimal $\hat{V}$. Ideally, $\hat{V}$ optimises the QFI and hardware efficiency. The compilation of such $\hat{V}$s is undergoing active research [40, 33].
A related opportunity concerns decoherence. Consider performing echo metrology using spins subject to spontaneous decay. The protocol may achieve the decayed probe state’s QFI [93]. Echo metrology merits extensions to sensors subject to general decoherence [58]. Perhaps we can gain insights from weak-value amplification and ICO metrology, which offer robustness with respect to certain noise (Sections IV and VI).
Beyond the parameter estimation of Sec. II, the echo-metrology template can facilitate further quantum information processing. The proof in Sec. III B generalises to quantum state discrimination and noise estimation [169]; the Dolinar receiver illustrates how [45]. Such a receiver optimally distinguishes two coherent states by implementing a displace-probe-replace protocol. Also, prepare-evolve-reverse sequences enable Hamiltonian amplification [12, 24, 163] and the detection of higher-order out-of-time-order correlators [2, 180]. More extensions might await discovery.
@@ IV Weak-value amplification
Weak measurements allow one to probe a target quantum system without disturbing it significantly [88]; von Neumann’s measurement model describes the process [88, 168]. Consider a target associated to a discrete Hilbert space. This target will entangle with the probe, unlike the classical-field target in echo metrology. Denote by $\hat{A}=\sum_{n}a_{n}\lvert a_{n}\rangle\!\langle a_{n}\rvert$ any observable defined on the target’s Hilbert space. Let $a_{\rm min}$ denote the least eigenvalue; and $a_{\rm max}$, the greatest.
@@
This $\hat{U}_{\alpha}$ differs from that in Sec. II (and so from the $\hat{U}_{\alpha}$ in echo metrology and some time-loop metrology) by acting on the target-and-probe composite, rather than on the probe alone. We then measure the probe’s position, $\hat{Q}$, projectively. If the probe begins with a wide wave function, $\Delta\phi\gg\alpha(a_{\max}-a_{\min})$, the interaction perturbs the joint state weakly. Across many trials, we infer the probability distribution over the possible final probe positions. The distribution peaks at $\alpha\langle\hat{A}\rangle\equiv\alpha\langle\Psi_{\mathrm{i}}\rvert\hat{A}\lvert\Psi_{\mathrm{i}}\rangle$.
@@ V Time-loop metrology
Agnostic and positronium sensing qualify not only as time-loop metrology, but also as echo metrology (Sec. III). The initial state $\lvert 0\rangle$, in step 1 of Sec. III B, is a product $\lvert 0,0\rangle$ of computational-basis states here. The preparatory unitary $\hat{V}$ there maps $\lvert 0,0\rangle$ to $\lvert\Psi^{-}\rangle$ here. The singlet-or-not measurement here follows from implementing $\hat{V}^{\dagger}$, then measuring whether the state is $\lvert 0,0\rangle$ (step 3 of Sec. III).
In the second adaptation of agnostic sensing, we introduce entanglement not in the preparation and measurement steps, but in the evolution [92]. An ancilla controls whether a single-qubit probe undergoes $\hat{U}_{\alpha}$ or $\hat{U}_{\alpha}^{\dagger}$. We prepare the ancilla in a superposition, such that the probe undergoes a superposition of the evolutions. This strategy mirrors ICO metrology (Sec. VI), relying on quantum control. However, we do not apply the unitaries with indefinite order here. To infer $\alpha$, one measures the ancilla. This protocol can achieve the optimal FI accessible with a single-qubit probe. However, one might not be able to invert the unknown $\hat{U}_{\alpha}$, or coherently control it, easily.
@@
Finally, we can elaborate on the relationship between time-loop metrology and echo metrology. Agnostic and positronium metrology qualify as echo metrology while involving CTC simulations. Do other echo-metrology experiments involve CTCs or variations thereon? Some echo-metrology experiments involve bosons, multiparticle entanglement, or submaximal entanglement. In contrast, researchers have argued that qubits, bipartite entanglement, and maximal entanglement can simulate CTCs [111]. Can we generalise the theory of CTC simulations to accommodate the latter setups? The effort may illuminate not only connections amongst classes of time-reverse metrology, but also the foundations of CTC theory.

### v2

@@ Abstract
Quantum metrology involves the application of quantum resources to enhance measurements. Several communities have developed quantum-metrology strategies that leverage effective time reversals. These strategies, we posit, form four classes. First, mirror metrology begins with a preparatory unitary and ends with that unitary’s time-reverse. The protocol amplifies the visibility of a small parameter to be sensed. Similarly, weak-value amplification enhances a weak coupling’s detectability. The technique exhibits counterintuitive properties captured by a retrocausal model. Using the third strategy, one simulates closed timelike curves, worldlines that loop back on themselves in time. The fourth strategy involves indefinite causal order, which characterises channels applied in a superposition of orderings. We review these four strategies, which we unify under the heading of time-reverse metrology. We also outline opportunities for this toolkit in quantum metrology; quantum information science; quantum foundations; atomic, molecular, and optical physics; and solid-state physics.
@@ I Introduction
    We can prepare a sensor with a unitary operation $\hat{V}$, couple the sensor to a target system, and then subject the sensor to $\hat{V}^{\dagger}$. The $\hat{V}$ and $\hat{V}^{\dagger}$ magnify the target’s impact on the sensor state. We call this strategy mirror metrology.
@@
Mirror Weak-value Time-loop ICO metrology
@@
Mirror In $\leftarrow$, a unitary realises Some mirror- In $\leftarrow$, we invert a
@@ II Technical introduction
Often, when performing parameter estimation, we infer the parameter $\alpha\in\mathbb{R}$ in a unitary $\hat{U}_{\alpha}\coloneqq e^{-i\alpha\hat{H}}$. The generator $\hat{H}$ is a Hermitian operator. Some target system produces $\hat{U}_{\alpha}\,$; example targets include an external field, as well as a physical system that couples to the probe. We aim to infer $\alpha$ with the least error possible, given our resources. [^2] In some time-loop metrology (Sec. V) and ICO metrology (Sec. VI), the inferred parameter characterises a nonunitary channel. We illustrate parameter estimation here with $\hat{U}_{\alpha}$ for simplicity and due to $U_{\alpha}$’s role in mirror metrology (Sec. III), weak-value amplification (Sec. IV), some time-loop metrology, and some ICO metrology.
@@
III Mirror metrology
Mirror metrology involves a unitary $\hat{V}$ and the time-reverse $\hat{V}^{\dagger}$ thereof. We adapt the name from mirror benchmarking, which features the same elements [21_Mayer_Theory]. Section III A motivates the protocol intuitively. In Sec. III B, we detail the protocol and prove that it achieves the information-bearing probe state’s QFI. Section III C illustrates the protocol with examples that involve squeezing. Section III D describes opportunities for future work.
III A Intuition behind mirror metrology
The following story helps motivate mirror metrology, although without explaining its behaviour rigorously. Consider a classical chaotic system, such as a double pendulum. It undergoes the protocol sketched in Fig. 1: the system begins at some point in phase space, then evolves under its chaotic Hamiltonian. A small perturbation nudges the system’s phase-space point. Then, the system evolves backward in time under the Hamiltonian. The chaos magnifies the typical such perturbation, by definition.
Figure 1: Motivation for mirror metrology: Sketch of a classical chaotic system’s trajectory through phase space. The system evolves forward under its Hamiltonian, undergoes a perturbation, and evolves backward under its Hamiltonian. The chaos magnifies the perturbation, rendering it easier to detect.
This protocol follows a pattern similar to that of mirror metrology. The perturbation above plays the role of the $e^{-i\alpha\hat{H}}$ introduced in Sec. II. Forward and reverse evolutions enhance the signal’s detectability. References [101, 103, 1] suggested the analogy. However, mirror metrology need not involve chaos; it relies on nontrivial, if not necessarily chaotic, evolution.
III B Introduction to, and analysis of, mirror metrology
Below, we detail mirror metrology. We then prove that it achieves the QFI of the information-bearing probe state. Finally, we review literature about mirror metrology (that does not involve squeezing, covered in Sec. III C).
Mirror metrology features the following setup and protocol. Consider a quantum probe whose Hilbert space contains a simple (unentangled, unsqueezed, etc.) state $\lvert 0\rangle$. A Hamiltonian $\hat{H}$ generates a unitary $e^{-i\alpha\hat{H}}$ parameterised by an $\alpha\in\mathbb{R}$ of magnitude $|\alpha|\ll 1$. The following protocol implements mirror metrology:
@@ III B Introduction to, and analysis of, mirror metrology
The final equality follows from step 1 of the protocol. This FI equals the QFI of $e^{-i\alpha\hat{H}}\,|\psi\rangle$ at $\mathcal{O}(\alpha)$, by Eq. (5). Thus, mirror metrology achieves the QFI.
Previous works state or implicitly leverage this result. Reference [114] synopsises the argument above. In [33], a similar argument focusses on certain $\hat{V}$s and $\hat{H}$s. Other references address general mirror metrology, arguing that it achieves the QFI of $e^{-i\alpha\hat{H}}\hat{V}\lvert 0\rangle$ [104, 175]. The review [3] illustrates mirror metrology with examples. Researchers have also extended mirror metrology to include ancilla systems [33].
Experiments have realised mirror metrology. The protocol has been performed (in the absence of squeezing) with trapped ions [99, 100, 116], a Rydberg atom and an oscillator [132], and nitrogen-vacancy centres in diamond [146]. Many optical interferometers implement mirror metrology, casting $V$ as a beamsplitter unitary [26, 144].
III C Mirror metrology based on squeezing
In much mirror metrology, the preparation involves squeezing. Squeezing is an operation that reduces one observable’s uncertainty at the expense of another observable’s [76]. We illustrate squeezing-based mirror metrology with single-mode squeezing, spin squeezing, and two-mode squeezing.
First, we motivate single-mode mirror metrology. Consider a bosonic mode – equivalently, a harmonic oscillator – associated to a position operator $\hat{Q}$ and a momentum operator $\hat{P}$. Suppose that a field imprints a parameter $\alpha$ on the oscillator’s state at a rate $g$, through the Hamiltonian $\hat{H}=g\alpha\hat{Q}$. This Hamiltonian kicks the oscillator’s momentum: $e^{i\hat{H}t}\,\hat{P}\,e^{-i\hat{H}t}=\hat{P}-g\alpha t\hat{\mathbbm{1}}$. The momentum’s displacement reveals $\alpha$, if we know $g$. Suppose we have prepared the oscillator in a state $\lvert\psi\rangle$. The momentum displacement’s expectation value forms an estimator:
@@ III C Mirror metrology based on squeezing
Squeezing facilitates mirror metrology for sensing $\alpha$. The probe begins in $\lvert 0\rangle$, undergoes squeezing by $e^{-ir\hat{H}_{\rm sq}(\varphi)}$, and evolves under $\hat{H}$. We then antisqueeze the sensor (evolve it under $e^{ir\hat{H}_{\rm sq}(\varphi)}$). Finally, we measure $\{\lvert 0\rangle\!\langle 0\rvert,\mathbbm{1}-\lvert 0\rangle\!\langle 0\rvert\}$. The unitaries condense into [^6] We apply the identity $e^{ir\hat{H}_{\rm sq}(\varphi{=}\pi/2)}\,\hat{Q}\,e^{-ir\hat{H}_{\rm sq}(\varphi{=}\pi/2)}=e^{r}\,\hat{Q}$.
@@
Not only a bosonic mode, but also a collection of spins, can undergo squeezing. [^7] The spins are typically qubits. However, we use the term spins for generality and for consistency with the literature. Various Hamiltonians enact spin squeezing, as discussed in App. A. Examples include the one-axis-twisting [93], the two-axis-twisting [93], the Lipkin-Meshkov-Glick [105], and general spin–spin interacting [19] Hamiltonians. Mirror metrology offers two benefits when implemented with spin squeezing. First, an estimator’s variance can scale in Heisenberg fashion with the number of spins [40]. Second, the strategy offers robustness with respect to readout error [40, 123, 96]. Neutral-atom experiments have showcased these advantages [78, 37, 103].
An early squeezing-based mirror-metrology proposal involves two-mode squeezing [178]. The protocol features two bosonic modes, associated to annihilation operators $\hat{a}$ and $\hat{b}$. Suppose that mode $a$ can undergo the phase shift $e^{i\alpha\hat{a}^{\dagger}\hat{a}}$. To sense $\alpha$, we prepare both modes in their vacuum states: $\lvert 0,0\rangle$. The unitary $\hat{S}(r)\coloneqq e^{r(\hat{a}^{\dagger}\hat{b}^{\dagger}-\hat{b}\hat{a})}$ introduces two-mode squeezing. Mode $a$ undergoes the phase shift. The unitary $\hat{S}^{-1}(r)$ unsqueezes the modes. Finally, we measure each mode’s photon number. This protocol achieves the QFI of $e^{i\alpha\hat{a}^{\dagger}\hat{a}}\,\hat{S}(r)\lvert 0,0\rangle$.
@@
III D Outlook for mirror metrology
The mirror-metrology template offers opportunities in and beyond the parameter estimation of Sec. II. Below, we present open questions about implementations. We then summarise possible applications to quantum information processing outside parameter estimation. Section V details how certain mirror-metrology experiments simulate CTCs.
The mirror-metrology protocol begins with a unitary $\hat{V}$ (step 1 in Sec. III B). What, generally, is the optimal form of $\hat{V}$? Consider any $\hat{V}$ that prepares an equal-weight superposition $(\lvert h_{\rm min}\rangle+e^{-i\phi}\lvert h_{\rm max}\rangle)/\sqrt{2}\,$, wherein $\phi\in[0,2\pi)$. Every such $\hat{V}$ can provide the QFI, maximised over $\hat{V}^{\prime}$, of $e^{-i\alpha\hat{H}}\hat{V}^{\prime}\lvert 0\rangle$ [Eq. (5)]. Yet implementing such $\hat{V}$s may pose challenges; for example, they may cost many computational gates. Furthermore, noise can affect the optimal $\hat{V}$. Ideally, $\hat{V}$ optimises the QFI and hardware efficiency. The compilation of such $\hat{V}$s is undergoing active research [40, 33].
A related opportunity concerns decoherence. Consider performing mirror metrology using spins subject to spontaneous decay. The protocol may achieve the decayed probe state’s QFI [92]. Mirror metrology merits extensions to sensors subject to general decoherence [58]. Perhaps we can gain insights from weak-value amplification and ICO metrology, which offer robustness with respect to certain noise (Sections IV and VI).
Beyond the parameter estimation of Sec. II, the mirror-metrology template can facilitate further quantum information processing. The proof in Sec. III B generalises to quantum state discrimination and noise estimation [168]; the Dolinar receiver illustrates how [45]. Such a receiver optimally distinguishes two coherent states by implementing a displace-probe-replace protocol. Also, prepare-evolve-reverse sequences enable Hamiltonian amplification [12, 24, 162] and the detection of higher-order out-of-time-order correlators [2, 179]. More extensions might await discovery.
@@ IV Weak-value amplification
Weak measurements allow one to probe a target quantum system without disturbing it significantly [87]; von Neumann’s measurement model describes the process [87, 167]. Consider a target associated to a discrete Hilbert space. This target will entangle with the probe, unlike the classical-field target in mirror metrology. Denote by $\hat{A}=\sum_{n}a_{n}\lvert a_{n}\rangle\!\langle a_{n}\rvert$ any observable defined on the target’s Hilbert space. Let $a_{\rm min}$ denote the least eigenvalue; and $a_{\rm max}$, the greatest.
@@
This $\hat{U}_{\alpha}$ differs from that in Sec. II (and so from the $\hat{U}_{\alpha}$ in mirror metrology and some time-loop metrology) by acting on the target-and-probe composite, rather than on the probe alone. We then measure the probe’s position, $\hat{Q}$, projectively. If the probe begins with a wide wave function, $\Delta\phi\gg\alpha(a_{\max}-a_{\min})$, the interaction perturbs the joint state weakly. Across many trials, we infer the probability distribution over the possible final probe positions. The distribution peaks at $\alpha\langle\hat{A}\rangle\equiv\alpha\langle\Psi_{\mathrm{i}}\rvert\hat{A}\lvert\Psi_{\mathrm{i}}\rangle$.
@@ V Time-loop metrology
Agnostic and positronium sensing qualify not only as time-loop metrology, but also as mirror metrology (Sec. III). The initial state $\lvert 0\rangle$, in step 1 of Sec. III B, is a product $\lvert 0,0\rangle$ of computational-basis states here. The preparatory unitary $\hat{V}$ there maps $\lvert 0,0\rangle$ to $\lvert\Psi^{-}\rangle$ here. The singlet-or-not measurement here follows from implementing $\hat{V}^{\dagger}$, then measuring whether the state is $\lvert 0,0\rangle$ (step 3 of Sec. III).
In the second adaptation of agnostic sensing, we introduce entanglement not in the preparation and measurement steps, but in the evolution [91]. An ancilla controls whether a single-qubit probe undergoes $\hat{U}_{\alpha}$ or $\hat{U}_{\alpha}^{\dagger}$. We prepare the ancilla in a superposition, such that the probe undergoes a superposition of the evolutions. This strategy echos ICO metrology (Sec. VI), relying on quantum control. However, we do not apply the unitaries with indefinite order here. To infer $\alpha$, one measures the ancilla. This protocol can achieve the optimal FI accessible with a single-qubit probe. However, one might not be able to invert the unknown $\hat{U}_{\alpha}$, or coherently control it, easily.
@@
Finally, we can elaborate on the relationship between time-loop metrology and mirror metrology. Agnostic and positronium metrology qualify as mirror metrology while involving CTC simulations. Do other mirror-metrology experiments involve CTCs or variations thereon? Some mirror-metrology experiments involve bosons, multiparticle entanglement, or submaximal entanglement. In contrast, researchers have argued that qubits, bipartite entanglement, and maximal entanglement can simulate CTCs [110]. Can we generalise the theory of CTC simulations to accommodate the latter setups? The effort may illuminate not only connections amongst classes of time-reverse metrology, but also the foundations of CTC theory.